Angles G and H are complementary. If measurement of angle G =3x + 6 and measurement of angle H = 2x - 11, what are the measure of each angle?
Answer by Mathtut(3670)

3x+6+2x-11=90
:
5x=95
:
x=19
:
3(19)+6=63
2(19)-11=27